THE INFRARED SPECTRUM OF THE NH_{4}^{+} CATION TRAPPED IN SOLID NEON

WARREN E. THOMPSON and \underline{MARILYN E. JACOX}, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

Both of the infrared-active vibrational fundamentals of the NH_{4}^{+} cation trapped in solid neon have been observed near the corresponding gas-phase band centers. Experimental conditions favoring the stabilization of this species–the first protonated molecule to be prepared in a neon matrix environment–will be explored. The study was extended to the partially and fully deuterated ammonium cations, for which the first infrared spectroscopic data will be reported.